# SpikeCLIP: A Contrastive Language-Image Pretrained Spiking Neural Network

### v3

@@ Abstract
Spiking Neural Networks (SNNs) have emerged as a promising alternative to conventional Artificial Neural Networks (ANNs), demonstrating comparable performance in both visual and linguistic tasks while offering the advantage of improved energy efficiency. Despite these advancements, the integration of linguistic and visual features into a unified representation through spike trains poses a significant challenge, and the application of SNNs to multimodal scenarios remains largely unexplored. This paper presents SpikeCLIP, a novel framework designed to bridge the modality gap in spike-based computation. Our approach employs a two-step recipe: an “alignment pre-training” to align features across modalities, followed by a “dual-loss fine-tuning” to refine the model’s performance. Extensive experiments reveal that SNNs achieve results on par with ANNs while substantially reducing energy consumption across various datasets commonly used for multimodal model evaluation. Furthermore, SpikeCLIP maintains robust image classification capabilities, even when dealing with classes that fall outside predefined categories. This study marks a significant advancement in the development of energy-efficient and biologically plausible multimodal learning systems.
@@ 1 Introduction
Artificial Neural Networks (ANNs) equipped with advanced deep learning techniques have demonstrated remarkable performance across a broad spectrum of visual and language tasks, sometimes even surpassing human capabilities (Krizhevsky et al., 2017; Graves and Jaitly, 2014; Mikolov et al., 2013). However, the significant computational power and energy required to operate these cutting-edge deep neural models have been steadily escalating over the past decade. This substantial energy expenditure poses a major barrier to the wide-spread application of deep learning. In contrast to ANNs, Spiking Neural Networks (SNNs) utilize discrete spikes for computation and information transmission, mirroring the energy efficiency of biological neurons. Neuromorphic hardware based on spike computation is now available and offers a more energy-efficient solution for implementing deep neural networks compared to specialized hardware such as GPUs. It has been reported that improvements in energy consumption of up to $2\backsim 3$ orders of magnitude when compared to conventional ANN acceleration on embedded hardware (Azghadi et al., 2020; Ceolini et al., 2020; Davies et al., 2021). Therefore, SNNs offer a promising computing paradigm to deal with large volumes of data using spike trains for information representation in a more energy-efficient manner.
@@
Many neuromorphic systems now allow us to simulate software-trained models without performance loss. However, mature on-chip training solutions are not yet available. SNNs are typically trained using software simulation platforms, and then the resultant models are uploaded onto neuromorphic hardware for inference. Unlike their conventional ANN counterparts, it remains a great challenge to train SNNs due to the non-differentiability of discrete spikes, a challenge that persists even within software simulation environments. Nevertheless, the performance gap between SNNs and ANNs has been constantly narrowing through intensive research conducted on SNNs in recent years, and can even vanish on some vision tasks (Cao et al., 2015; Diehl et al., 2015; Rueckauer et al., 2017; Hu et al., 2018; Yin et al., 2020; Fang et al., 2021; Zhou et al., 2023a, b). Although relatively fewer studies have explored the effectiveness of SNNs in Natural Language Processing (NLP) (Diehl et al., 2016; Rao et al., 2022), recent work indicates that spiking convolution networks can achieve results comparable to those of ANNs in multiple language datasets with significantly lower energy consumption (Lv et al., 2022). Despite these advances, existing research has predominantly focused on single-modality tasks, and the potential for applying SNNs to multimodal contexts remains largely unexplored.
Figure 1: An illustration of a two-step recipe for training multimodal SNNs. Initially, we pre-train SpikeCLIP by distilling knowledge from the conventional CLIP. A readout layer is introduced to facilitate knowledge distillation from traditional ANNs into SNNs by interpreting SNN states and mapping them to feature representations. Subsequently, we fine-tune the pre-trained SpikeCLIP on downstream datasets. In addition to the standard cross-entropy loss, we introduce a regularization term based on the Kullback-Leibler divergence. This term penalizes significant discrepancies between the feature representations generated by SpikeCLIP and those produced by CLIP, thereby preserving the generalization capacity obtained during pre-training and enhancing zero-shot learning capabilities.
@@
To extend the application of SNNs to multimodal contexts, a significant challenge lies in the alignment of features extracted from multimodal inputs in the form of spike trains. If such an alignment or a mapping can be achieved, we can translate texts and images into their representations of meaning, and these representations can be leveraged to assess meaning similarity across different modalities. Moreover, if this mapping can make predictions about concept categories to unseen inputs, it would suggest that the underlying model has successfully captured meaning representations across modalities using discrete spikes. The approach to meaning representation that currently dominates the field of machine learning relies on distributed semantic representations, also known as embeddings. One straightforward method for converting the spike trains produced by a collection of neurons into a representation is to interpret the firing rates as the activations of the representation (Cao et al., 2015). With this conversion, cross-modal mapping can be achieved through contrastive learning, as demonstrated in the dual-stream CLIP (Radford et al., 2021).
@@
Obtaining a high-performing cross-modal mapping through contrastive learning requires a substantial quantity of text-image pairs for the joint training of an image encoder and a text encoder. It has been reported that as many as $400$ million pairs were used to train the CLIP. Regrettably, these $400$ million text-image pairs are not yet publicly accessible. Fortunately, the pre-trained CLIP has been made available to the research community, which enables us to employ the Knowledge Distillation (KD) technique (Hinton et al., 2015) to train our spiking variant, which we have named SpikeCLIP. However, distilling knowledge from ANNs and transferring it to SNNs is non-trivial, as there is no simple solution for representing negative values in SNNs. To circumvent this problem, we used a readout layer to interpret the states of an SNN and map them to the feature representations that are amenable to knowledge distillation. This approach draws on the principles of liquid state machines (Maass and Markram, 2004; Maass, 2011), a subclass of reservoir computer (Röhm and Lüdge, 2018; Tanaka et al., 2019) that uses SNNs for dynamic data processing.
@@
After pre-training SpikeCLIP through the distillation of knowledge from the conventional CLIP, we proceed to finetune it on downstream datasets for image classification tasks. To enhance performance on instances whose classes fall outside predefined categories, we employ a dual-loss strategy that minimizes the cross-entropy between the predicted probabilities and actual distributions, while introducing a regularization term based on the Kullback-Leibler (KL) divergence. The purpose of this term is to impose penalties on any significant discrepancies between the feature representations generated by the SNNs and those produced by the CLIP model, which helps to maintain the generalization capacity gained during the pre-training stage. Our ablation study demonstrated that such regularization can significantly enhance the performance of SpikeCLIP, particularly for images with categories not presented in the labels of a given dataset. To overcome the non-differentiable issue, we generalized the backpropagation algorithm with surrogate gradients (Zenke and Vogels, 2021) to train the SNNs. The contribution of this study can be summarized as follows:
  • We have demonstrated that the features extracted from multimodal inputs (i.e., texts and images) can be effectively aligned in the form of spike trains, which can then be leveraged to predict concept categories for unseen inputs.
@@
  • We introduce a two-step method for training multimodal SNNs, involving pre-training for cross-modal alignment through knowledge distillation, followed by dual-loss finetuning with surrogate gradients (see Figure 1).
  • This study is among the first to demonstrate that well-trained SNNs can match the performance of their ANN counterparts in multimodal classification tasks and even exhibit zero-shot learning capabilities.
@@ 2 Related Work
Spike neural networks have drawn considerable attention in recent years due to their potential to realize artificial intelligence while greatly reducing energy consumption. Several training methods have been proposed to mitigate the non-differentiable of SNNs, which can generally be categorized into conversion-based and spike-based methods. Conversion-based methods are to train a non-spiking network first and convert it into an SNN that inherits the learned weights of the non-spiking network (Cao et al., 2015; Diehl et al., 2015; Sengupta et al., 2019; Rathi et al., 2020; Lv et al., 2022). The advantage of such methods is that the non-differentiability of discrete spikes can be circumvented and the burden of training in the temporal domain is partially removed. On the other hand, spike-based methods train SNNs using spike-timing information in either a supervised or unsupervised manner. The majority of research in this line relies on the gradients estimated by a differentiable approximate function so that gradient descent can be applied with backpropagation using spike times (Bohte et al., 2002; Booij and tat Nguyen, 2005) or backpropagation using spikes (i.e., backpropagation through time) (Shrestha and Orchard, 2018; Hunsberger and Eliasmith, 2015; Bellec et al., 2018; Huh and Sejnowski, 2018). These two training methods, or their combinations, have been investigated to train SNNs for single-modality tasks including computer vision or language processing.
@@
Many studies have shown that SNNs can yield competitive results in vision (mostly classification) tasks (Cao et al., 2015; Diehl et al., 2015; Rueckauer et al., 2017; Shrestha and Orchard, 2018; Sengupta et al., 2019). Cao et al. (2015) pioneered a method that successfully converted a deep convolutional neural network into an SNN by interpreting the activations as firing rates. To minimize performance degradation during the conversion process, Diehl et al. (2015) introduced a novel weight normalization method to regulate firing rates, which enhances the performance of SNNs in image classification tasks without additional training time. Sengupta et al. (2019) pushed SNNs to go deeper by investigating residual architectures and introducing a layer-by-layer weight normalization method. To mitigate performance loss after the conversion, Bu et al. (2023) suggested to use of a quantization clip-floor-shift activation function instead of the ReLU function in ANNs so that the spiking patterns of SNNs could be more accurately simulated during the training of the corresponding ANNs. Inspired by the well-established transformer architecture (Vaswani et al., 2017), Zhou et al. (2022) introduced a spiking version called Spikeformer, which was further refined to reduce reliance on floating-point computations (Zhou et al., 2023a, b). By leveraging transformer-like architectures, Spikeformer and its variants achieved state-of-the-art results in image classification tasks.
@@
While the application of SNNs in the field of computer vision has been extensively investigated, their effectiveness in NLP tasks has been relatively less explored (Diehl et al., 2016; Rao et al., 2022; Lv et al., 2022). Rao et al. (2022) demonstrated that long-short-term memory (LSTM) units could be implemented on spike-based neuromorphic hardware using the spike frequency adaptation mechanism. Diehl et al. (2016) used pre-trained word embeddings in their TrueNorth implementation of a recurrent neural network and achieved $74\%$ accuracy in a question classification task. However, an external projection layer is required to project word embeddings to the vectors with positive values that can be further converted into spike trains. Lv et al. (2022) proposed a two-step recipe of “conversion + fine-tuning” to train spiking neural networks for NLP. Initially, a normally trained ANN is converted into an SNN by duplicating its architecture and weights. Subsequently, the converted SNN undergoes fine-tuning. They showed that SNNs trained using this method can yield competitive results on both English and Chinese datasets compared to their ANN counterparts. Lv et al. (2023) extended the Spikeformer to make it possible to process language tasks, resulting in the SpikeBERT. The SpikeBERT not only outperformed state-of-the-art SNNs but also achieved results comparable to BERTs on certain text classification datasets while significantly reducing energy consumption.
@@
Although the extensive exploration of SNNs in single-modality tasks, their potential application in multimodal contexts remains largely untapped. However, deep neural networks have shown their efficacy in multimodal modeling, which can be broadly categorized into single-stream architectures such as OSCAR (Li et al., 2020) and SimVLM (Wang et al., 2021), and dual-stream architectures like CLIP and WenLan (Huo et al., 2021). Single-stream models process multimodal inputs through a unified architecture, where all types of data are combined at an early stage before being fed into task-specific layers while dual-stream models process each modality through separate sub-networks first and then merge the outputs of these networks at a later stage. Single-stream models are less flexible in handling the specific characteristics of each modality, as every type of input is treated uniformly through the same network layers. In contrast, dual-stream models can optimize the processing for each modality independently, which can lead to better handling of the unique features of each data type. Our preliminary attempts with single-stream architectures have not yielded the desired results probably because it is hard for SNNs to capture interactions between different types of data early in the forms of spike trains. However, we found that SNNs, when implemented with a dual-stream architecture and trained with a method that combines alignment pre-training with dual-loss fine-tuning, can rival the performance of their ANN counterparts in various multimodal classification tasks. Moreover, they display the capacity for zero-shot learning.
@@ 3 Method
Figure 2: Owing to the challenges associated with early-stage integration of different modalities by spiking signals, we opted for dual-stream architectures for multimodal modeling. This choice allows for more tailored processing of each modality before integration. The spiking image encoder is built upon the Spikingformer (Zhou et al., 2023a), a spiking variant of the Transformer, which employs a spiking self-attention mechanism and spiking MLPs as replacements for the original components. Although other architectures could potentially be applied, we designed a simple spiking multilayer perceptron to construct the text encoder of SpikeCLIP. A set of integrate-and-fire neurons is used to convert data from any modalities into spike trains for subsequent processing within SNNs.
To facilitate the application of SNNs in image-text multimodal tasks, it is critical to first align the semantic features extracted from both texts and images in the form of spike trains. Given that single-stream models exhibit less flexibility in accommodating the specific characteristics of each modality and the challenges of integrating different modalities at early processing stages by SNNs, we chose to use dual-stream architectures for multimodal modeling with SNNs. This approach allows for more tailored processing of each modality before integration. The paradigm of pre-training models on a large dataset to learn general features, followed by fine-tuning on task-specific datasets, has demonstrated considerable success in NLP and computer vision. Following this training paradigm, we first pre-train SNNs by distilling the knowledge from the CLIP, and then the pre-trained networks are fine-tuned on task-specific datasets with a dual-loss function. A readout layer is used to interpret the states of an SNN, which overcomes the barrier posed by the discrepancy between float-valued feature representations in conventional ANNs and temporal spiking representations used in SNNs.

### v4

@@ Abstract
Spiking Neural Networks (SNNs) have emerged as a promising alternative to conventional Artificial Neural Networks (ANNs), demonstrating comparable performance in both visual and linguistic tasks while offering the advantage of improved energy efficiency. Despite these advancements, the integration of linguistic and visual features into a unified representation through spike trains poses a significant challenge, and the application of SNNs to multimodal scenarios remains largely unexplored. This paper presents SpikeCLIP, a novel framework designed to bridge the modality gap in spike-based computation. Our approach employs a two-step recipe: an “alignment pre-training” to align features across modalities, followed by a “dual-loss fine-tuning” to refine the model’s performance. Extensive experiments reveal that SNNs achieve results on par with ANNs while substantially reducing energy consumption across various datasets commonly used for multimodal model evaluation. Furthermore, SpikeCLIP maintains robust image classification capabilities, even when dealing with classes that fall outside predefined categories. This study marks a significant advancement in the development of energy-efficient and biologically plausible multimodal learning systems. Our code is available at https://github.com/Lvchangze/SpikeCLIP.
keywords:
Spiking Neural Networks , Multimodal Models
journal: Neural Networks
[…]
[1]organization=School of Computer Science, Fudan University, city=Shanghai, postcode=200433, country=China
[…]
[2]organization=University College London, city=London, country=UK
{graphicalabstract}
Figure 1: The training framework of SpikeCLIP.
Figure 2: The architecture of SpikeCLIP.
{highlights}
Spiking-Based Multimodal Feature Alignment: This work is among the first to demonstrate that multimodal features extracted from text and images can be effectively aligned using spike train representations. These aligned representations enable zero-shot prediction of concept categories in previously unseen inputs.
Novel Training Algorithm for Multimodal SNNs: We propose a two-step method for training multimodal SNNs, which includes pre-training for cross-modal alignment via knowledge distillation, followed by dual-loss fine-tuning with surrogate gradients.
Comprehensive Experimental Evaluation: We conduct extensive experiments to assess the performance of SpikeCLIP on image classification tasks. Additionally, we perform ablation studies to demonstrate the model’s zero-shot capability and its reduction in energy consumption.
@@ 1 Introduction
Artificial Neural Networks (ANNs) equipped with advanced deep learning techniques have demonstrated remarkable performance across a broad spectrum of visual and language tasks, sometimes even surpassing human capabilities [1, 2, 3]. However, the significant computational power and energy required to operate these cutting-edge deep neural models have been steadily escalating over the past decade. This substantial energy expenditure poses a major barrier to the widespread application of deep learning. In contrast to ANNs, Spiking Neural Networks (SNNs) utilize discrete spikes for computation and information transmission, mirroring the energy efficiency of biological neurons. Neuromorphic hardware based on spike computation is now available and offers a more energy-efficient solution for implementing deep neural networks compared to specialized hardware such as GPUs. It has been reported that improvements in energy consumption of up to $2\backsim 3$ orders of magnitude when compared to conventional ANN acceleration on embedded hardware [4, 5, 6]. Therefore, SNNs offer a promising computing paradigm to deal with large volumes of data using spike trains for information representation in a more energy-efficient manner.
@@
In reality, many neuromorphic systems experience performance loss during migration from simulation to hardware due to quantization or varying hardware support for operations, as shown in [7]. However, mature on-chip training solutions are not yet available. SNNs are typically trained using software simulation platforms, and then the resultant models are uploaded onto neuromorphic hardware for inference. Unlike their conventional ANN counterparts, it remains a great challenge to train SNNs due to the non-differentiability of discrete spikes, a challenge that persists even within software simulation environments. Recent intensive research [8, 9, 10, 11, 12, 13, 14, 15] on SNNs has significantly narrowed the performance gap between SNNs and ANNs, with this gap even disappearing in some vision tasks. Although relatively fewer studies have explored the effectiveness of SNNs in Natural Language Processing (NLP) [16, 17], recent work indicates that spiking convolution networks can achieve results comparable to those of ANNs in multiple language datasets with significantly lower energy consumption [18]. Despite these advances, existing research has predominantly focused on single-modality tasks, and the potential for applying SNNs to multimodal contexts remains largely unexplored.
Figure 3: An illustration of our two-step training approach for multimodal SNNs. First, we pre-train SpikeCLIP by distilling knowledge from conventional CLIP, using a readout layer to map SNN states to floating-point feature representations. Then, we fine-tune SpikeCLIP on downstream datasets, adding a regularization term based on Kullback-Leibler divergence on the training loss.
@@
A key challenge in extending the application of SNNs to multimodal contexts is the alignment of features extracted from multimodal inputs into spike train representations. Achieving this alignment or mapping would enable the translation of texts and images into meaningful representations, which can then be used to assess the similarity of meaning across different modalities. Moreover, if this mapping can make predictions about concept categories to unseen inputs, it would suggest that the underlying model has successfully captured meaning representations across modalities using discrete spikes. The approach to meaning representation that currently dominates the field of machine learning relies on distributed semantic representations, also known as embeddings. One straightforward method for converting the spike trains produced by a collection of neurons into a representation is to interpret the firing rates as the activations of the representation [8]. With this conversion, cross-modal mapping can be achieved through contrastive learning, as demonstrated in the dual-stream CLIP [19].
@@
Obtaining a high-performing cross-modal mapping through contrastive learning requires a substantial quantity of text-image pairs for the joint training of an image encoder and a text encoder. It has been reported that as many as $400$ million pairs were used to train the CLIP. Regrettably, these $400$ million text-image pairs are not yet publicly accessible. Fortunately, the pre-trained CLIP has been made available to the research community, which enables us to employ the Knowledge Distillation (KD) technique [20] to train our spiking variant, which we have named SpikeCLIP. However, distilling knowledge from ANNs and transferring it to SNNs is non-trivial, as there is no simple solution for representing negative values in SNNs. To circumvent this problem, we used a readout layer to interpret the states of an SNN and map them to the feature representations that are amenable to knowledge distillation. This approach draws on the principles of liquid state machines [21, 22], a subclass of reservoir computer [23, 24] that uses SNNs for dynamic data processing.
@@
As shown in Figure 3), after pre-training SpikeCLIP through the distillation of knowledge from the conventional CLIP, we proceed to fine-tune it on downstream datasets for image classification tasks. To enhance performance on instances whose classes fall outside predefined categories, we employ a dual-loss strategy that minimizes the cross-entropy between the predicted probabilities and actual distributions, while introducing a regularization term based on the Kullback-Leibler (KL) divergence. The purpose of this term is to impose penalties on any significant discrepancies between the feature representations generated by the SNNs and those produced by the CLIP model, which helps to maintain the generalization capacity gained during the pre-training stage. Our ablation study demonstrated that such regularization can significantly enhance the performance of SpikeCLIP, particularly for images with categories not presented in the labels of a given dataset. To overcome the non-differentiable issue, we generalized the backpropagation algorithm with surrogate gradients [25] to train the SNNs.
The contribution of this study can be summarized as follows:
  • 1.
    Spiking-Based Multimodal Feature Alignment: This work is among the first to demonstrate that multimodal features extracted from text and images can be effectively aligned using spike train representations. These aligned representations enable zero-shot prediction of concept categories in previously unseen inputs.
  • 2.
@@
    Novel Training Algorithm for Multimodal SNNs: We propose a two-step method for training multimodal SNNs, which includes pre-training for cross-modal alignment via knowledge distillation, followed by dual-loss fine-tuning with surrogate gradients.
  • 3.
    Comprehensive Experimental Evaluation: We conduct extensive experiments to assess the performance of SpikeCLIP on image classification tasks. Additionally, we perform ablation studies to demonstrate the model’s zero-shot capability and its reduction in energy consumption.
@@ 2.1 Training Methods for SNNs
Spike neural networks have drawn considerable attention in recent years due to their potential to realize artificial intelligence while greatly reducing energy consumption. Several training methods have been proposed to mitigate the non-differentiable of SNNs, which can generally be categorized into conversion-based and spike-based methods. Conversion-based methods are to train a non-spiking network first and convert it into an SNN that inherits the learned weights of the non-spiking network [8, 9, 26, 27, 18]. The advantage of such methods is that the non-differentiability of discrete spikes can be circumvented and the burden of training in the temporal domain is partially removed. On the other hand, spike-based methods train SNNs using spike-timing information in either a supervised or unsupervised manner.
The majority of research [28, 29, 30, 31] in this line relies on the surrogate gradients training method, which estimated the back gradients with a differentiable approximate function so that gradient descent can be applied with backpropagation using spike times [32, 33] or backpropagation using spikes (i.e., backpropagation through time). Gu et al. [34], Ma et al. [35], Wang et al. [36] explore complementary aspects of SNNs, which contribute to a deeper understanding of spatiotemporal credit assignment, the role of noise in computation and learning, and adaptive gradient smoothing techniques. These two training methods, or their combinations, have been investigated to train SNNs for single-modality tasks including computer vision or language processing. In this paper, we propose a novel two-stage training method for multi-modal SNNs based on surrogate gradients.
2.2 SNNs for Single-Modal Tasks
@@
Many studies have shown that SNNs can yield competitive results in vision (mostly classification) tasks [8, 9, 10, 29, 26]. Cao et al. [8] pioneered a method that successfully converted a deep convolutional neural network into an SNN by interpreting the activations as firing rates. To minimize performance degradation during the conversion process, Diehl et al. [9] introduced a novel weight normalization method to regulate firing rates, which enhances the performance of SNNs in image classification tasks without additional training time. Sengupta et al. [26] pushed SNNs to go deeper by investigating residual architectures and introducing a layer-by-layer weight normalization method. To mitigate performance loss after the conversion, Bu et al. [37] suggested using a quantization clip-floor-shift activation function instead of the ReLU function in ANNs so that the spiking patterns of SNNs could be more accurately simulated during the training of the corresponding ANNs. Inspired by the well-established Transformer architecture [38], Zhou et al. [39] introduced a spiking version called Spikeformer, which was further refined to reduce reliance on floating-point computations [14, 15]. By leveraging Transformer-like architectures, Spikeformer and its variants achieved state-of-the-art results in image classification tasks.
@@
While the application of SNNs in the field of computer vision has been extensively investigated, their effectiveness in NLP tasks has been relatively less explored [16, 17, 18]. Rao et al. [17] demonstrated that long-short-term memory (LSTM) units could be implemented on spike-based neuromorphic hardware using the spike frequency adaptation mechanism. Diehl et al. [16] used pre-trained word embeddings in their TrueNorth implementation of a recurrent neural network and achieved $74\%$ accuracy in a question classification task. However, an external projection layer is required to project word embeddings to the vectors with positive values that can be further converted into spike trains. Lv et al. [18] proposed a two-step recipe of “conversion + fine-tuning” to train spiking neural networks for NLP. Initially, a normally trained ANN is converted into an SNN by duplicating its architecture and weights. Subsequently, the converted SNN undergoes fine-tuning. They showed that SNNs trained using this method can yield competitive results on both English and Chinese datasets compared to their ANN counterparts. Lv et al. [40], Bal and Sengupta [41] extended the Spiking Transformer [39] to make it possible to process language tasks, resulting in the SpikeBERT. Their SNNs not only outperformed state-of-the-art models but also achieved results comparable to BERTs on certain text classification datasets while significantly reducing energy consumption.
2.3 SNNs for Multi-Modal Tasks
@@
Deep neural networks have shown their efficacy in multimodal modeling, which can be broadly categorized into single-stream architectures such as OSCAR [42] and SimVLM [43], and dual-stream architectures like CLIP and WenLan [44]. Single-stream models process multimodal inputs through a unified architecture, where all types of data are combined at an early stage before being fed into task-specific layers while dual-stream models process each modality through separate sub-networks first and then merge the outputs of these networks at a later stage. Single-stream models are less flexible in handling the specific characteristics of each modality, as every type of input is treated uniformly through the same network layers. In contrast, dual-stream models can optimize the processing for each modality independently, which can lead to better handling of the unique features of each data type. Our preliminary attempts with single-stream architectures have not yielded the desired results probably because it is hard for SNNs to capture interactions between different types of data early in the forms of spike trains. However, we found that SNNs, when implemented with a dual-stream architecture and trained with a method that combines alignment pre-training with dual-loss fine-tuning, can rival the performance of their ANN counterparts in various multimodal classification tasks. Moreover, they display the capacity for zero-shot learning.
Although the extensive exploration of SNNs in single-modality tasks, their potential application in multimodal contexts remains largely untapped. Several recent studies [45, 46, 47] have demonstrated the feasibility of integrating multimodal information into SNN models. However, these studies primarily focus on modalities such as speech and images, with few addressing the fusion of text and image modalities. Panchev [48] proposes a simple SNN module for robots to better understand language instructions; however, this module does not qualify as a multimodal SNN since it excludes the integration of both text and image data. In contrast, our work is among the first to focus on the fusion of text and image modalities. We introduce a two-stage training method for multimodal SNNs and achieve promising performance across several benchmark datasets.
@@ 3 Method
Figure 4: Overview of the architecture of SpikeCLIP. We use dual-stream architectures for multimodal modeling. The spiking image encoder is based on Spikingformer [14], while a simple spiking MLP is designed for the text encoder of SpikeCLIP, with integrate-and-fire neurons converting data into spike trains for SNN processing.
3.1 Challenges and Motivations
To enable the use of SNNs in image-text multimodal tasks, it is essential to align the semantic features extracted from both texts and images into spike train representations. Given the limited flexibility of single-stream models in accommodating the unique characteristics of each modality, and the challenges of early-stage modality integration in SNNs, we adopt a dual-stream architecture for multimodal modeling. This approach allows for more specialized processing of each modality before integration.
The successful paradigm of pre-training on large datasets to learn general features, followed by fine-tuning on task-specific datasets, has shown strong results in NLP and computer vision. In line with this, we first pre-train SNNs by distilling knowledge from CLIP, then fine-tune the pre-trained networks on task-specific datasets using a dual-loss function. We utilize a readout layer (see Section 3.5) to interpret the hidden states of the SNN, overcoming the challenge posed by the discrepancy between the floating-point feature representations of conventional ANNs and the temporal spiking representations of SNNs.